OpenStudy (anonymous):

find the exact trigonometric ratios for the angle whose radian measure is 4pi/3

OpenStudy (campbell_st):

well this is a 3rd quadrant angle where tan is positive, sin and cos are both negative in the 3rd quadrant, the angles are \[\pi + x\] then you know \[\frac{4\pi}{3} = \pi + x\] find the value of x, the you need to find tan(x), -sin(x) and -cos(x) they will all be exact values

OpenStudy (campbell_st):

to get the values draw a triangle